# SoK: LLM-based Log Parsing

### v1

@@ Abstract.
Log data, generated by software systems, provides crucial insights for tasks like monitoring, root cause analysis, and anomaly detection. Due to the vast volume of logs, automated log parsing is essential to transform semi-structured log messages into structured representations. Traditional log parsing techniques often require manual configurations, such as defining log formats or labeling data, which limits scalability and usability. Recent advances in large language models (LLMs) have introduced the new research field of LLM-based log parsing, offering potential improvements in automation and adaptability. Despite promising results, there is no structured overview of these approaches since this is a relatively new research field with the earliest advances published in late 2023. This paper systematically reviews $29$ LLM-based log parsing methods, comparing their capabilities, limitations, and reliance on manual effort. We analyze the learning and prompt-engineering paradigms employed, efficiency- and effectiveness-enhancing techniques, and the role of LLMs in the parsing process. We aggregate the results of the survey in a large table comprising the characterizing features of LLM-based log parsing approaches and derive the general process of LLM-based log parsing, incorporating all reviewed approaches in a single flow chart. Additionally, we benchmark seven open-source LLM-based log parsers on public datasets and critically assess their reproducibility. Our findings summarize the advances of this new research field and provide insights for researchers and practitioners seeking efficient and user-friendly log parsing solutions, with all code and results made publicly available for transparency.
parsing, log data, large language models
copyright: nonedoi: XXXXXXX.XXXXXXXjournalvolume: Xjournalnumber: Xarticle: Xpublicationmonth: X
@@ 1. Introduction
Log data consists of recorded information generated by logging statements in software applications, systems, devices, or networks during operation. It provides a chronological record of events and contains valuable insights for various tasks, such as monitoring (Chen et al., 2019; Wang et al., 2021), root cause analysis (He et al., 2021; Notaro et al., 2023), and anomaly detection (Landauer et al., 2023a; He et al., 2016a). Given that modern computer systems can generate millions of log lines per hour, manual processing is infeasible, necessitating fast and scalable automated approaches. For instance, in 2013 the Alibaba Cloud system was reported to produce approximately 100 to 200 million log lines per hour (Mi et al., 2013). However, extracting meaningful information from raw log data requires transforming it into a structured format. Since logs are typically semi-structured, following general patterns but containing variable log messages, direct analysis is challenging. To enable log analytics tasks like anomaly detection, logs must first be parsed into a structured representation, making log parsing a critical step in log analysis (Wurzenberger et al., 2019).
@@
Log parsing refers to the extraction of structured data from semi-structured logs. Parsing the logs into the structured format can simply be done by regular expression matching if the correct log templates are given. Most methods therefore formulate log parsing as a log template extraction problem (He et al., 2017; Jiang et al., 2024a). Ground-truth templates are defined somewhere in logging statements in the source code and are therefore often not available. A wide range of log parsing techniques (Shima, 2016; Jiang et al., 2008; Du and Li, 2016; Sedki et al., 2022; Le and Zhang, 2023b) have therefore been proposed to overcome this issue.
@@
Many state of the art log parsing methods require manual configuration or other manual actions at some point in the parsing process. Approaches may involve labeling of some of the logs from the training set (Jiang et al., 2024a; Xu et al., 2024; Le and Zhang, 2023b), definition of log formats (Li et al., 2023) or regular expressions for preprocessing (Yu et al., 2023; He et al., 2017; Sedki et al., 2022), or other parameters (Yu et al., 2023) that fit the parsing algorithm to the data. From a user’s perspective the configuration of parsing algorithms might seem overwhelming since it requires expertise or an analysis of the data at hand. The reason log parsing methods require manual actions is that humans have both semantic and syntactic understanding (Jiang et al., 2024a), but more importantly, a natural generalization capability and potentially expert knowledge about logs and log parsing. Many log parsers have already achieved semantic (Le and Zhang, 2023b; Liu et al., 2022; Jiang et al., 2024a) and syntactic abilities (He et al., 2017; Dai et al., 2022; Yu et al., 2023). While syntax-based parsers focus on characteristics like occurrence frequencies, or word or log length, to determine the constant part of logs, semantic-based parsers leverage the semantic differences between static and variable parts of log messages (Jiang et al., 2024a; Xiao et al., 2024). Knowledge about logs and how to parse them can be learned by language models, such as BERT (Devlin et al., 2019) or its descendants, by fine-tuning them with logs and their templates (Le and Zhang, 2023b; Tao et al., 2022). However, these methods often lack generalization and have to be trained or fine-tuned with labeled logs, preferably with logs of the exact same log type as the target logs (Ma et al., 2024a). Generalization arises from pre-training on a large corpus of diverse datasets, allowing models to perform well also on novel tasks and possibly enhanced through fine-tuning (FT) (Gao et al., 2021) or in-context learning (ICL) (Brown et al., 2020). This is where large language models (LLMs) come into play.
@@
With the emergence of LLMs the new research field of LLM-based log parsing arised. In late 2023, Le and Zhang (Le and Zhang, 2023a) reported notable performance of ChatGPT [^1] https://openai.com/index/chatgpt/; accessed 10-March-2025 in a naive setting, where the LLM was presented individual logs and simply asked for their templates. In the meantime, other approaches adopted LLMs for this purpose as well, building sophisticated frameworks that utilize LLMs in various ways to either parse logs directly or enhance log parsing, by adopting learning paradigms for LLMs such as ICL (Brown et al., 2020) or fine-tuning (Gao et al., 2021). Given the fact that the number of unique templates is significantly lower than the number of logs and the average rate of newly discovered templates decreases as more logs are processed (Jiang et al., 2024a), some LLM-based parsers (Jiang et al., 2024a) are even able to achieve runtime performance comparable with the fast and well-known parser Drain (He et al., 2017).
@@
To the best of our knowledge, there is an absence of a structured overview and an objective comparison of papers and methods concerning LLM-based log parsing, as this is a rather novel research field which only recently emerged from the popular research field of generative LLMs. So far there is merely a survey paper about LLM-based log analysis techniques by Akhtar et al. (Akhtar et al., 2025), which provides a rudimentary overview on using LLMs for log parsing and the techniques employed. Consequently, this work aims to undertake a systematic review of the literature to identify the commonalities and variabilities of the existing approaches and to present our gained insights. Our main focus is an overview of LLM-based log parsing with emphasis on the users’ perspective — thus, how the LLM is applied and how much manual effort the approaches require. The aim is therefore to provide a comprehensive overview of the existing approaches and techniques, enabling anyone interested in (LLM-based) log parsing to quickly and easily identify the most suitable solution for their specific use case.
The aforementioned goals are formulated as the following research questions:
@@
  • RQ4: What experiment design choices influence the comparability of the evaluation results between LLM-based log parsing approaches?
@@
We summarize the main contributions of the present work as follows:
  • We provide a systematic overview of the existing methods, based on a set of feature definitions that characterize LLM-based log parsing approaches, or the corresponding papers, respectively.
@@
  • Based on the literature review and our evaluation results, we answer $5$ research questions and derive further findings that may benefit the future research and design of log parsers.
  • Finally, we make the source code and all the results of our evaluation publicly available in our GitHub repository [^2] https://github.com/ait-aecid/LLM-log-parsing.git to ensure transparency and reproducibility.
The remainder is structured as follows: Section 2 describes the background and the core concepts and how we understand them. Section 3 outlines the survey method, including the search strategy and the definition of the reviewed features. Section 4 presents the results of the literature review in a large table and describes the findings of each feature. Section 5 explains the evaluation setup, while Sec. 6 presents the evaluation results. We discuss the findings from literature and evaluation in Sec. 7 and conclude our paper in Sec. 8.
2. Background and Related Work
@@ 2.1. Log Parsing
Log parsing is the foremost step in log analytics and denotes the process of extracting structured data from logs. Logs are given as single- or multi-line events and are available in textual form (Landauer et al., 2023a). The diverse and often un- or semi-structured nature of logs makes log parsing a difficult task. It is the precondition for many log analytics tasks, hence the effectiveness of downstream tasks is therefore inherently dependent on the effectiveness of the parser and also determines the runtime efficiency of the whole process, which is especially important for online (real-time) or incremental approaches, such as live-monitoring for cybersecurity (Jiang et al., 2024a). Incremental approaches are designed to handle continuous data streams, updating the parsing model dynamically as new log data arrives. This allows for real-time analysis and quicker responses to emerging issues.
In this work, we specifically focus on template extraction, since the extraction of the variables from the log messages is trivial once the (correct) templates are identified, for example, with regular expression matching. The template extraction process is therefore the key component in log parsing.
@@ 2.2. Log Templates
A log template is a predefined format used to log system events in software applications. It typically includes placeholders for essential details like timestamps, log levels (e.g., INFO, ERROR), message content, contextual metadata and more (Li et al., 2023). Previous research differs between the unstructured log message part of a log and the structured log headers (He et al., 2016b). The structured part is thereby straightforward to extract with regular expressions and it has been shown that parsing only the log message part improves the parsing accuracy (He et al., 2016b). Many methods (He et al., 2017; Du and Li, 2016; Xu et al., 2024; Du and Li, 2016), therefore require the log format of a dataset as input parameters. An example of a log template from the Apache dataset (Zhu et al., 2023) is given in Fig. 1. The log format parameter for this example would be “[<Time>] [<Level>] <Content>”, whereas “Time” and “Level” refers to log headers and “Content” refers to the log message. In the remainder, we understand “log format” as this parameter defining the positions of the log headers and of the content within different log types, or different datasets, respectively.
@@ 2.3. Large Language Models (LLMs)
Large Language Models (LLMs) have evolved significantly, transitioning from early statistical and rule-based approaches to deep learning architectures powered by Generative Pretrained Transformers (GPTs). These models leverage self-attention mechanisms and large-scale datasets to generate human-like text, enabling advancements in natural language processing (NLP) tasks such as summarization, translation, content generation, and complex problem-solving. A key enhancement in LLMs is instruction tuning which specializes the pretrained LLMs to follow instructions provided in prompts. Instruction tuning enables effective ICL. Instead of modifying model parameters, ICL allows LLMs to learn from prompts that include instructions, examples, and queries, making them more adaptable to specialized applications (Zhang et al., 2024b).
The term LLM is only loosely defined, hence, for this work, we define LLMs as generative pretrained transformers (GPTs) such as ChatGPT, Mistral or Qwen. Models like BERT (Devlin et al., 2019) (encoder-only) are thereby excluded, as we are only interested in typically larger, generative models that excel in both understanding and generating text, making them suitable for a broader range of tasks. Many works included BERT into their definition of LLMs which complicated the manual filtering process after the keyword search.
@@ 2.4. Log Parsing with LLMs
Large-scale model series like ChatGPT and many more have demonstrated potential in this domain. Some approaches fine-tune smaller LLMs (Mehrabi et al., 2024) or BERT (Le and Zhang, 2023b) to this specific task while reporting superior parsing performance to state of the art approaches. However, fine-tuning LLMs for log parsing is resource-intensive in terms of computational costs, runtime, as well as labeled examples, making ICL a compelling alternative (Xiao et al., 2024). Research shows that ICL enhances LLMs’ performance in logical reasoning and fact retrieval, suggesting its viability for structured log analysis (Brown et al., 2020).
Despite these advantages, integrating LLMs into log parsing presents challenges. These models are not inherently designed for log parsing, leading to inconsistencies in token classification and errors in log templates. Such inaccuracies can negatively impact downstream tasks like anomaly detection (Zhu et al., 2019). Additionally, given the vast volume of real-world log data, efficient processing is crucial. The computational demands of LLMs contribute to inference overhead and network latency, necessitating optimizations for practical deployment (Jiang et al., 2024a).

### v2

# System Log Parsing with Large Language Models: A Review

@@ Abstract.
Log data provides crucial insights for tasks like monitoring, root cause analysis, and anomaly detection. Due to the vast volume of logs, automated log parsing is essential to transform semi-structured log messages into structured representations. Recent advances in large language models (LLMs) have introduced the new research field of LLM-based log parsing. Despite promising results, there is no structured overview of the approaches in this relatively new research field with the earliest advances published in late 2023. This work systematically reviews 29 LLM-based log parsing methods. We benchmark seven of them on public datasets and critically assess their comparability and the reproducibility of their reported results. Our findings summarize the advances of this new research field, with insights on how to report results, which data sets, metrics and which terminology to use, and which inconsistencies to avoid, with code and results made publicly available for transparency.
log parsing, log data, large language models, LLMs
copyright: nonejournalyear: nonacmdoi: XXXXXXX.XXXXXXXjournalvolume: Xjournalnumber: Xarticle: Xpublicationmonth: Xccs: Computing methodologies Information extractionccs: Software and its engineering Software performance
@@ 1. Introduction
Log data consists of recorded information generated by logging facilities in software applications, systems, devices, or networks during operation. It provides a chronological record of events and contains valuable insights for various tasks, such as monitoring (Chen et al., 2019; Wang et al., 2021), root cause analysis (He et al., 2021; Notaro et al., 2023), and anomaly detection (Landauer et al., 2023a; He et al., 2016a). Given that modern computer systems can generate millions of log lines per hour, manual processing is infeasible, necessitating fast and scalable automated approaches. For instance, in 2013 the Alibaba Cloud system was reported to produce approximately 100 to 200 million log lines per hour (Mi et al., 2013). However, extracting meaningful information from raw log data requires transforming it into a structured format. Since logs are typically semi-structured, following general patterns but containing variable log messages, direct analysis is challenging. To enable log analytics tasks like anomaly detection, logs must first be parsed into a structured representation, making log parsing a critical step in log analysis (Wurzenberger et al., 2019).
@@
Log parsing refers to the extraction of structured data from semi-structured logs. Parsing the logs into the structured format can simply be achieved by regular expression matching if the correct log templates are given. Most methods therefore formulate log parsing as a log template extraction problem (He et al., 2017; Jiang et al., 2024a; Vaarandi and Bahsi, 2024; Yu et al., 2023; Sedki et al., 2022; Dai et al., 2022). Ground-truth templates are defined somewhere in logging statements in the source code and are therefore often not available. A wide range of log parsing techniques (Shima, 2016; Jiang et al., 2008; Du and Li, 2016; Sedki et al., 2022; Le and Zhang, 2023b) have therefore been proposed to overcome this issue.
@@
Many state-of-the-art log parsing methods require manual configuration or other manual actions at some point in the parsing process. Approaches may involve labeling of some of the logs from the training set (Jiang et al., 2024a; Xu et al., 2024; Le and Zhang, 2023b), definition of log formats (Li et al., 2023) or regular expressions for preprocessing (Yu et al., 2023; He et al., 2017; Sedki et al., 2022), or other parameters (Yu et al., 2023) that fit the parsing algorithm to the data. From a user’s perspective the configuration of parsing algorithms might seem overwhelming since it requires expertise or an analysis of the data at hand. The reason log parsing methods require manual actions is that humans have both semantic and syntactic understanding (Jiang et al., 2024a), but more importantly, a natural generalization capability and potentially expert knowledge about logs and log parsing. Many log parsers have already achieved semantic (Le and Zhang, 2023b; Liu et al., 2022) and syntactic abilities (He et al., 2017; Dai et al., 2022; Yu et al., 2023), but many do not achieve satisfactory performance on data other than LogHub (Zhu et al., 2023) as reported by (Fu et al., 2022; Zhang et al., 2024a; Li et al., 2024a) or require substantial amounts of labeled logs for training their models (Xiao et al., 2024). While syntax-based parsers focus on characteristics like occurrence frequencies, or word or log length, to determine the static parts of logs, semantic-based parsers leverage the semantic differences between static and variable parts of log messages (Jiang et al., 2024a; Xiao et al., 2024). Knowledge about logs and how to parse them can be learned by language models, such as BERT (Devlin et al., 2019) or its descendants, by fine-tuning them with logs and their templates (Le and Zhang, 2023b; Tao et al., 2022; Liu et al., 2022). However, these methods often lack generalization and have to be trained or fine-tuned (Gao et al., 2021) with labeled logs, preferably with logs of the exact same log type as the target logs (Ma et al., 2024a). Generalization arises from pretraining on a large corpus of diverse datasets, allowing models to perform well also on novel tasks and possibly enhanced through fine-tuning (Gao et al., 2021) or in-context learning (ICL) (Brown et al., 2020). This is where large language models (LLMs) come into play.
@@
With the emergence of LLMs the new research field of LLM-based log parsing arised. In late 2023, Le and Zhang (Le and Zhang, 2023a) reported notable performance of ChatGPT [^1] https://openai.com/index/chatgpt/; accessed 10-March-2025 in a naive setting, where the LLM was presented individual logs and simply asked for their templates. In the meantime, other approaches adopted LLMs for this purpose as well, building sophisticated frameworks that utilize LLMs in various ways to enhance log parsing by adopting learning paradigms for LLMs such as ICL (Brown et al., 2020) or fine-tuning (Gao et al., 2021). Given the fact that the number of unique templates is significantly lower than the number of logs and the average rate of newly discovered templates decreases as more logs are processed (Jiang et al., 2024a), some LLM-based parsers (Jiang et al., 2024a; Pei et al., 2024) are even able to achieve runtime performance comparable with the fast and well-known parser Drain (He et al., 2017).
@@
To the best of our knowledge, a structured overview and an objective comparison of papers and methods concerning LLM-based log parsing is currently lacking, as this is a rather novel research field which only recently emerged from the popular research field of generative LLMs. So far there is only a preprint survey paper by Akhtar et al. (Akhtar et al., 2025), which provides an overview of using LLMs for log analysis, including log parsing. Although their discussion on log parsing is concise, it lays a valuable foundation for further innovation and development in this area. Consequently, our work aims to undertake a systematic review of the literature, identifying the commonalities and variabilities of the existing approaches, defining common terminologies and reporting schemes, highlighting performance enhancing paradigms, and deriving recommendations for future advancements in this field. Our main focus is on investigating LLM-based log parsing methods with emphasis on researchers seeking to apply or develop novel log parsing techniques. Thus, additional emphasis is placed on the application of the LLM and the manual effort required for the implementation of the various approaches.
The aforementioned goals are approached by answering the following research questions:
@@
  • RQ4: Which experiment design choices hinder the comparability, and hence which guidelines should be adopted in terms of configuration, datasets used, evaluation metrics, and reporting to make results comparable?
@@
We summarize the main contributions of this work as follows:
  • We provide a systematic overview of the existing methods and techniques based on a set of feature definitions that characterize LLM-based log parsing approaches, or the corresponding papers, respectively.
@@
  • Based on the literature review and our evaluation results, we answer $5$ research questions, deriving recommendations for future research and the design of (LLM-based) log parsers.
  • Finally, we make the source code and all the results of our evaluation publicly available in our GitHub repository (Beck, 2025) to ensure transparency and reproducibility.
The remainder is structured as follows: Section 2 describes the background and the core concepts. Section 3 describes the related work. Section 4 outlines the survey method, including the literature search strategy and the definition of the reviewed features. Section 5 presents the results of the literature review in a large table and describes the findings of each feature. Section 6 explains the evaluation setup, while Sec. 7 presents the evaluation results. We discuss the findings from literature and evaluation in Sec. 8, and conclude our paper in Sec. 9.
2. Background
@@ 2.1. Log Parsing
Log parsing is the foremost step in log analytics and denotes the extraction of structured data from logs. Logs are given as single- or multi-line events, available in textual form (Landauer et al., 2023a). The diverse and un- or semi-structured nature of logs makes log parsing a difficult task. It is the precondition for many log analytics tasks, hence the effectiveness of downstream tasks is inherently dependent on the effectiveness of the parser. Parsing also determines the runtime efficiency of the whole process, which is especially important for online or incremental approaches, such as live-monitoring for cybersecurity (Jiang et al., 2024a). Incremental approaches are designed to handle continuous data streams, updating the parsing model dynamically as new log data arrives, allowing for real-time analysis and quick responses to emerging issues.
In this work, we specifically focus on template extraction, since the extraction of the variables from the log messages is trivial once the (correct) templates are identified, for example, with regular expression matching. The template extraction process is therefore the key component of log parsing.
@@ 2.2. Log Templates
A log template is a predefined format used to log system events in software applications. It typically includes placeholders for essential details like timestamps, log levels (e.g., INFO, ERROR), message content, contextual metadata and more (Li et al., 2023). Previous research differentiates between the unstructured log message part of a log and the structured log headers (He et al., 2016b). The structured part is thereby straightforward to extract with regular expressions. It has been shown that parsing only the unstructured log message part improves the parsing accuracy for conventional (non-LLM-based) parsers (He et al., 2016b). Many methods (He et al., 2017; Du and Li, 2016; Xu et al., 2024; Du and Li, 2016), therefore require the log format ($\neq$ log template) of a dataset as an input parameter. An example of a log template from the Apache dataset (Zhu et al., 2023) is given in Fig. 1. The log format parameter for this example would be “[<Time>] [<Level>] <Content>”, whereas “Time” and “Level” refers to log headers and “Content” refers to the log message. In the remainder, we understand the “log format” as this parameter defining the positional indicators of the log headers and of the content within different log types, or different datasets, respectively.
@@ 2.3. Large Language Models (LLMs)
Large Language Models (LLMs) have evolved significantly, transitioning from early statistical and rule-based approaches to deep learning architectures powered by Generative Pretrained Transformers (GPTs). These models leverage self-attention mechanisms and large-scale datasets to generate human-like text, enabling advancements in natural language processing (NLP) tasks such as summarization, translation, content generation, and complex problem-solving. A key enhancement in LLMs is instruction tuning which specializes the pretrained LLMs to follow instructions provided in prompts. Instruction tuning enables effective in-context learning (ICL). Instead of modifying model parameters, ICL allows LLMs to learn from prompts that include instructions, examples, and queries, making them more adaptable to specialized applications (Zhang et al., 2024b).
The term LLM is only loosely defined, hence, for this work, we define LLMs as generative pretrained transformers (GPTs) such as ChatGPT, Mistral or Qwen. Models like BERT (Devlin et al., 2019) (encoder-only) are thereby excluded, as we are only interested in generative models that excel in both understanding and generating text, with increased generalization capabilities compared to BERT, making them suitable for a broader range of tasks.
@@ 2.4. Log Parsing with LLMs
Large-scale model series like ChatGPT and many more have demonstrated potential in this domain. Some approaches fine-tune smaller LLMs (Mehrabi et al., 2024) or BERT (Le and Zhang, 2023b) to this specific task while reporting superior parsing performance to state-of-the-art approaches. However, fine-tuning LLMs for log parsing is resource-intensive in terms of computational costs, runtime, and labeling examples, making ICL a compelling alternative (Xiao et al., 2024). Research shows that ICL enhances LLMs’ performance in logical reasoning and fact retrieval, suggesting its viability for structured log analysis (Brown et al., 2020). However, LLM inference is also costly. The computational demands of LLMs contribute to inference overhead and network latency, but also energy consumption and thus, an increased carbon footprint. Using LLMs for automation therefore necessitates optimizations for practical and energy efficient deployment (Jiang et al., 2024a; Xiao et al., 2024). Regardless of the efficiency, LLMs are not inherently designed for log parsing, leading to inconsistencies in token classification and errors in log templates. Inaccuracies in the parsed templates can negatively impact downstream tasks like anomaly detection (Zhu et al., 2019). Despite the aforementioned challenges there is a variety of approaches that address these concerns and provide solutions, making LLM-based log parsing a strong alternative to conventional log parsing.
3. Related Work
Existing surveys and benchmarks have significantly contributed to the understanding and evaluation of log parsing. These works are particularly relevant to our work, as they provide a comprehensive overview of the current state of the field, highlighting key methodologies, challenges, and advancements: The work of Zhu et al. (Zhu et al., 2019) provides a comprehensive benchmark of log parsing techniques, as well as LogHub (Zhu et al., 2023), which became the common dataset collection used in many log analysis evaluations. The labels of the LogHub datasets were corrected by Khan et al. (Khan et al., 2022) to a unified template style. Jiang et al. (Jiang et al., 2024b) conducted a large-scale evaluation on state-of-the-art log parsing techniques while also providing the large-scale dataset collection, LogHub-2.0. It contains several million annotated logs and has been widely adopted for the evaluation of log analysis techniques (Jiang et al., 2024a; Xiao et al., 2024; Karanjai et al., 2024; Huang et al., 2024; Zhong et al., 2024). The large number of logs in LogHub-2.0 facilitated comprehensive efficiency evaluations, which are becoming increasingly important due to the large number of logs generated by modern computer systems (Mi et al., 2013; Zhang et al., 2023). Similarly to our work, Zhang et al. (Zhang et al., 2023) classify existing parsing approaches by a set of features, such as the parsing mode or the required amount of preprocessing effort. The researchers contemplate the challenges of state-of-the-art log parsing techniques, more specifically, the scarcity of public datasets, limited generalizability caused by intrinsic limitations, a lack of automation, and insufficient efficiency regarding the implementation.
The integration of LLMs into log analysis methods has been explored in the recent survey paper by Akhtar et al. (Akhtar et al., 2025). It provides a comprehensive survey of LLM-based techniques for log analysis, highlighting their potential to enhance efficiency and accuracy in processing substantial volumes of log data. The authors discuss various applications, including anomaly detection, fault monitoring, root cause analysis, threat detection, but also provides an overview log parsing. The paper explores the use of fine-tuning, ICL, and retrieval-augmented generation (RAG) (Lewis et al., 2020) for performance improvements in log parsing. It also reviews existing research, identifying common challenges such as the high resource consumption of LLMs, the need for frequent updates, and security concerns with closed-source models.